# mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt

### v3

@@ Abstract
Software vulnerabilities pose significant risks to the security and integrity of software systems. Although prior studies have explored vulnerability detection using deep learning and pre-trained models, these approaches often fail to provide the detailed explanations necessary for developers to understand and remediate vulnerabilities effectively. The advent of large language models (LLMs) has introduced transformative potential due to their advanced generative capabilities and ability to comprehend complex contexts, offering new possibilities for addressing these challenges. In this paper, we propose LLMVulExp, an automated framework designed to specialize LLMs for the dual tasks of vulnerability detection and explanation. To address the challenges of acquiring high-quality annotated data and injecting domain-specific knowledge, LLMVulExp leverages prompt-based techniques for annotating vulnerability explanations and fine-tunes LLMs using instruction tuning with Low-Rank Adaptation (LoRA), enabling LLMVulExp to detect vulnerability types in code while generating detailed explanations, including the cause, location, and repair suggestions. Additionally, we employ a Chain-of-Thought (CoT) based key code extraction strategy to focus LLMs on analyzing vulnerability-prone code, further enhancing detection accuracy and explanatory depth. Our experimental results demonstrate that LLMVulExp achieves over a 90% F1 score on the SeVC dataset, effectively combining high detection accuracy with actionable and coherent explanations. This study highlights the feasibility of utilizing LLMs for real-world vulnerability detection and explanation tasks, providing critical insights into their adaptation and application in software security.
11footnotetext:  Co-first authors, equally contributed.22footnotetext:  Corresponding author.
@@ I Introduction
A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2]. These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation. The increasing complexity and interconnectedness of software systems introduce the great challenge of identifying and mitigating these vulnerabilities effectively.
Current vulnerability detection techniques mainly include pattern based and deep learning based approaches. Pattern based approaches [3, 4] generally rely on manually defined rules to detect vulnerabilities. Deep learning based approaches [1, 2, 5, 6, 7] train the models using existing vulnerability data and various code representation techniques. Despite these advancements, existing methods often fall short of providing detailed explanations of detected vulnerabilities. This lack of robust explanatory capabilities impedes a comprehensive understanding and effective mitigation of vulnerabilities when applied to real-world usage. It is important to propose new techniques that can detect software vulnerabilities and provide additional explanations. Figure 1 shows an example of the vulnerability detection result with and without an explanation which provides comprehensive information for understanding and fixing the vulnerabilities.
In the context of enhancing detection methods, the advent of Large Language Models (LLMs) [8, 9, 10] offers a promising avenue. With their advanced generative capabilities, LLMs have demonstrated significant potential across a wide range of applications, such as natural language processing and machine translation. These models can generate extensive textual content and provide contextually relevant information, making them well-suited for tasks that require detailed explanations. However, a considerable gap exists between the current capabilities of LLMs, particularly open-source models, and the specific requirements of vulnerability detection and explanation. Without domain-specific knowledge of vulnerabilities, LLMs struggle to effectively detect vulnerabilities and offer accurate explanations, which demand a deep understanding of code structures, security contexts, and the intricate interplay between various software components. To bridge this gap, it is essential to enhance LLMs with specialized capabilities for vulnerability detection and explanation. Fine-tuning open-source LLMs [11, 12, 13] on task-specific data presents an effective path toward achieving model specialization.
Figure 1: Illustration of Vulnerability Detection Explanation.
The main challenges in this specialization fine-tuning lie in two areas: (1) acquiring high-quality vulnerability explanation data [1, 14, 10, 15, 16], and (2) injecting vulnerability-related domain knowledge into LLMs [17]. The improvement of model specialization depends heavily on high-quality, domain-specific data. However, in the field of vulnerability detection, there is a significant lack of annotated explanation data. Additionally, annotating such data typically requires costly manual labeling by domain experts, making the collection of large-scale annotated datasets a major challenge for model fine-tuning. Additionally, due to the inherent complexity of vulnerability detection and explanation tasks, merely providing task descriptions in prompts is insufficient for generating accurate answers. The model requires additional, critical vulnerability domain knowledge to identify and explain vulnerabilities accurately.
To tackle these challenges, this study introduces LLMVulExp, an innovative automated framework designed to specialize open-source LLMs for software vulnerability tasks. Our framework directly addresses the two key challenges identified: the acquisition of high-quality annotated explanatory data and the effective injection of domain-specific vulnerability knowledge into LLMs. To address the first challenge, we leverage prompt-based techniques to automatically annotate explanatory information for vulnerability data. This annotated data includes both the vulnerability location and a detailed explanation, providing a rich dataset for fine-tuning. This approach reduces reliance on costly manual labeling by domain experts, enabling the collection of large-scale, annotated datasets necessary for model specialization. In addition, we propose new evaluation metrics for vulnerability explanations and an automated ’LLM-as-a-judge’ evaluation method [18, 19], ensuring a robust and scalable approach to verify the correctness and coherence of the generated explanations. For the second challenge, we apply instruction tuning [20] with Low-Rank Adaptation (LoRA) fine-tuning methods [12] to refine the open-source LLMs for injecting sufficient vulnerability patterns and knowledge. By focusing on task-specific adaptation, we guide the model to develop a deeper understanding of vulnerability semantics. To further enhance the model’s structural understanding of vulnerabilities, we adopt a Chain-of-Thought (CoT) strategy [21], guiding the model to concentrate on key code snippets most prone to vulnerabilities.
@@
Overall, the contributions of this paper are threefold:
  • We introduce a comprehensive and effective workflow for training, inferring, and evaluating LLMs specifically on vulnerability detection and explanation tasks.
  • We pioneer the exploration of the effectiveness of LLM-based vulnerability explanation generation and propose a tailored evaluation method and metrics for these explanations.
  • We conduct extensive experiments to evaluate and analyze the feasibility of LLM-based vulnerability explanations, providing valuable insights for the practical application of specialized models in vulnerability detection.
@@ II-A Vulnerability Detection
Prior studies proposed a series of deep learning approaches [1, 2, 5, 6, 7] to detect vulnerabilities. These methods have employed labeled vulnerability data to train neural networks, enabling the models to capture semantic features associated with vulnerabilities.
@@ II-B Vulnerability Explanation
Although substantial progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these methods for vulnerability explanation remains a significant challenge. Despite the critical importance of reducing reliance on security experts and aiding developers in timely and thorough vulnerability mitigation, this area remains underexplored. Only a few studies focus on the explanatory capabilities of deep learning based models for vulnerabilities.
VulDeeLocator [29] enhances a Bi-LSTM detector with an inner multiplication layer to forecast vulnerable statements. IVDetect [2] and LineVul [30] utilize subgraphs or attention weights from trained detectors to identify vulnerabilities. VELVET [31] integrates graph-based and sequence-based neural networks to prioritize vulnerable statements. LineVD [32] treats statement-level vulnerability detection as a node classification task, combining graph neural networks with a transformer-based model on the Program Dependency Graph(PDG). VulTeller [33] focuses on control and taint flows to detect dependencies for localizing vulnerabilities. VulExplainer [17] uses graph neural networks to locate fine-grained vulnerability information.
While neural network-based models offer promising semantic understanding for vulnerability explanation, accurately detecting and explaining vulnerabilities using LLMs remains a substantial challenge for general-purpose code models. Our work explores the feasibility of fine-tuning specialized LLMs for both vulnerability detection and explanation, addressing a critical gap in the research literature.
@@ III Methodology
❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs requires large quantities of high-quality domain-specific data. In the context of vulnerability detection and explanation, it is crucial to ensure the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples for each type. In this paper, we conduct the study on two C/C++ function-level vulnerability datasets: (1) SeVC [34], which contains four core vulnerability types and over 50,000 vulnerable code snippets, and (2) DiverseVul [14], which covers 295 real open-source projects and 150 CWE types. We select C/C++ as the programming language due to its diverse range of vulnerability types and examples, as well as the substantial dependence of fine-tuning on the availability of large-scale data.
@@
Dataset Ori. Vul # Ann. Vul # Vul-Type # Eval. Setting
SeVC 56,395 40,491 4 Single/Multi-Type
DiverseVul 18,945 9,161 10 Multi-Type(CWE)
@@
DiverseVul is a comprehensive C/C++ vulnerability dataset that includes 18,945 vulnerable functions and 330,492 non-vulnerable functions, extracted from 7,514 commits, covering 150 CWEs. As the largest real-world C/C++ vulnerability dataset, DiverseVul is characterized by longer code snippets, a broader range of projects, more diverse vulnerability types, and lower label noise. It serves as a benchmark for evaluating the capability of LLMs in handling real-world scenarios with a wider variety of vulnerabilities. For our fine-tuning and evaluation target, we selected the top ten most frequent CWE types, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
To ensure data quality, we deduplicate the vulnerability samples using the SHA-256 [35] hash method and downsample the non-vulnerable samples at a 1:1 ratio to match the vulnerable samples, achieving a balanced dataset, which aims to reduce training overhead and mitigates potential model bias. The details of the datasets are shown in Table I. We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerable and non-vulnerable samples to conduct our experiments.
@@
We implement this annotation process using GPT-3.5 [37], accessed via the API provided by OpenAI [38] due to its balance of annotation cost and efficiency. As part of our experimental study, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively. This effort fills a significant gap in the availability of vulnerability explanation annotations, enabling more effective training and evaluation of vulnerability detection and explanation models. To validate the effectiveness of the synthesized data, we indirectly evaluated the performance of the fine-tuned model in vulnerability detection tasks across multiple scenarios. Additionally, we propose evaluation metrics for vulnerability explanation quality, along with both manual and automated evaluation methods, to directly assess the quality of the generated explanations.
@@ III-B Key Code Extraction Based on Chain-of-Thought (CoT)
To address these challenges and improve the capability of large models to analyze complex code structures, we propose a novel Chain-of-Thought (CoT) enhancement method that incorporates automated key code extraction. This approach enables the model to focus on the most relevant parts of the code, thereby improving accuracy and interpretability. Using the prompt template illustrated in Figure LABEL:fig:prompt_key, we employ LLMs to automatically extract key statements by analyzing the semantic context of the code and the specific vulnerability type. These extracted statements form a structured CoT, guiding the fine-tuned model through a step-by-step process for targeted vulnerability detection, location, and explanation. For example, in a buffer overflow scenario, key code may include array declarations, bounds-checking statements, or pointer dereferences that contribute directly to the vulnerability. By integrating these extracted key statements into the prompts during fine-tuning, the model can systematically reference this crucial information, enhancing both the interpretability and precision of the vulnerability explanation process.
@@ IV Results
  • RQ4: How does the key code extraction impact detection performance?
@@ IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
Accurate detection is the foundation for correct vulnerability explanation. In this research question (RQ), we first discuss the detection performance of the fine-tuned specialized vulnerability models across various scenarios.
@@
Dataset. As discussed in Section III, we select the SeVC [34] dataset and the DiverseVul dataset [14]. Based on the characteristics of the two datasets, we constructed three different vulnerability detection tasks to evaluate the detection performance of LLMVulExp: binary classification vulnerability detection (SeVC), coarse-grained multi-class vulnerability detection (SeVC), and CWE type-based multi-label vulnerability detection (DiverseVul).
Backbone LLMs. In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [39] as our primary model for the experiments discussed in this chapter. CodeLlama is initialized with the weights of Llama2 [8] and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities. Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability LLMs, we design three sub-research questions:
  • (RQ1.1) How effective is the model in detecting a single specific type of vulnerability?
  • (RQ1.2) How effective is the model in detecting multiple types of vulnerabilities simultaneously?
  • (RQ1.3) How does the model perform in real-world scenarios with fewer data and more types of vulnerabilities to detect?
To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics. For the binary classification scenario in RQ1.1, we directly use these three metrics. For the multi-class classification in RQ1.2, we use Weighted-F1 and Macro-F1. For the multi-label classification in RQ1.3, we use Micro-F1 and Macro-F1. To test detection accuracy, we selected CodeT5 [40] and CodeBERT [41] as baselines and fine-tuned them for classification tasks by adding a linear classification layer. These models are chosen for two primary reasons: (1) As pre-trained code models, CodeT5 and CodeBERT have demonstrated consistent and superior performance on prior vulnerability detection benchmarks [16, 14], making them reliable baselines. (2) Their inclusion allows us to assess whether fine-tuned LLMs offer an advantage over existing pre-trained code models in terms of detection accuracy. It is important to note that the focus of this study is not on developing the most optimal detection model but rather on evaluating the explanation capabilities of fine-tuned LLMs. Therefore, using these two baselines provides a meaningful comparison without detracting from the primary objective of assessing the interpretability of LLM-generated explanations.
Implementation Details. We implement our approach using the Transformers[13] and PEFT[42] libraries on the PyTorch platform. All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs, with a token length limit set to 2048. The models are trained using the AdamW optimizer for 3 epochs with a batch size of 2 and data precision set to float16 to optimize memory usage and computational efficiency. We set the learning rate to 0.0003 and weight decay to 0.01. For the LoRA configuration, we set target_modules to {q_proj, k_proj, v_proj, o_proj} for SeVC and {q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj} for DiverseVul. The LoRA rank is set to 16, the LoRA scaling factor to 16, and the dropout rate to 0.05. For fine-tuning CodeBERT and CodeT5, we use the AdamW optimizer with the learning rate of 1e-5 and the weight decay of 1e-2, and train the models for 3 epochs. All other parameters are set to their default values.
@@
IV-A4 RQ1.3: Identification of a greater variety of vulnerability types.
In real-world software development environments, security risks arise from a diverse range of project types and complex code structures, often involving a broader spectrum of vulnerability categories. To better reflect these real-world conditions, we conduct a multi-label classification task using the top 10 CWE categories in the real-world vulnerability dataset DiverseVul, where a subset of the code samples is annotated with multiple CWE labels. Specifically, the task description is modified to emphasize CWE-type vulnerability detection and explanation, with the addition of a ‘[CWE]’ tag to generate a list of CWE categories relevant to the target code. Furthermore, CWE descriptions are incorporated as part of the output explanation, aiding the model in comprehending the context and semantics of each CWE type and leveraging these descriptions for vulnerability analysis and identification.
TABLE V: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.3).
Metric CWE ID Overall (Micro-F1)
787 125 119 20 703 416 476 190 200 399 Ours CodeT5 CodeBERT
Precision 79.1% 70.9% 79.3% 68.1% 46.9% 82.1% 74.4% 80.4% 63.6% 88.2% 74.4% 70.8% 65.3%
Recall 83.3% 72.5% 75.8% 71.0% 53.6% 78.0% 56.9% 84.9% 75.0% 83.3% 75.2% 63.1% 61.9%
F1 Score 81.2% 71.7% 77.5% 69.5% 50.0% 80.0% 64.4% 82.6% 68.9% 85.7% 74.8% 67.0% 63.5%
Experimental Results. The experimental results are presented in Table V. The findings demonstrate that after fine-tuning, the specialized vulnerability models achieved high detection accuracy across the 10 CWE categories, highlighting their capability to capture the distinctive patterns associated with each CWE type. Unlike the multi-class task in SeVC, the multi-label classification in DiverseVul does not require the model to distinguish between mutually exclusive categories. Instead, each CWE type has a more granular and explicit definition, mitigating the category association challenges observed in SeVC and thereby yielding improved performance. These results validate the practicality of employing LLMs for vulnerability detection and explanation in real-world software development scenarios. The following example of a generated explanation further illustrates the model’s capacity to analyze and identify CWE-based vulnerabilities effectively.
1 Network::FilterStatus Context::onNetworkNewConnection() { 2 onCreate(root_context_id_); 3 if (!wasm_->onNewConnection_) { 4 return Network::FilterStatus::Continue; 5 } 6 if (wasm_->onNewConnection_(this, id_).u64_ == 0) { 7 return Network::FilterStatus::Continue; 8 } 9 return Network::FilterStatus::StopIteration; 10 } {mdframed} [style=mystyle] [label] This function is vulnerable.

### v4

# Towards Explainable Vulnerability Detection with Large Language Models

@@ Abstract
Automated vulnerability detection tools often lack the explainability needed to assist developer remediation. While Large Language Models (LLMs) are promising, creating explainable detectors faces two key challenges: acquiring high-quality explanation data at scale and instilling models with autonomous, end-to-end reasoning capabilities. To address these challenges, we propose LLMVulExp, an innovative “teacher-student” framework. Our core contribution is a pipeline that resolves the data bottleneck: we guide a powerful “teacher” LLM to generate a large-scale corpus of high-quality explanations for known vulnerabilities. This data is then used to efficiently train a “student” LLM, via instruction tuning and LoRA, to autonomously detect and explain vulnerabilities in unseen code. Our second key contribution is a novel evaluation protocol for explanations, featuring metrics for Accuracy, Clarity, and Actionability, alongside an automated “LLM-as-a-judge” method. This work thus presents a complete and validated end-to-end framework for building and evaluating specialized, explainable LLMs for software security.
Index Terms:
Vulnerability Detection, Vulnerability Explanation, Large Language Model, Instruction Tuning.
@@ I Introduction
Software vulnerabilities are flaws or weaknesses in a system that can be exploited by an attacker to perform unauthorized actions [1, 2, 3], potentially leading to severe consequences such as data breaches, financial losses, and reputational damage. The growing complexity and interconnectedness of modern software systems pose significant challenges to the timely and effective identification and mitigation of these vulnerabilities.
Vulnerability detection techniques are predominantly divided into pattern-based and deep learning-based approaches. Pattern-based methods [4, 5] rely on manually crafted rules, while deep learning-based approaches [1, 2, 6, 7, 8, 9, 10] leverage existing vulnerability datasets and various code representation techniques to train predictive models. Despite their advancements, a critical limitation of existing methods is their failure to provide detailed, human-understandable explanations for the vulnerabilities they detect. This explanatory gap severely impedes a developer’s ability to comprehend, validate, and effectively mitigate the identified risks in practice. Therefore, there is a pressing need for techniques that not only detect vulnerabilities but also provide insightful explanations, as illustrated in Figure 1. The figure contrasts a standard detection result with an enriched one, highlighting how comprehensive explanations aid in understanding and remediation.
Figure 1: Illustration of Vulnerability Detection with and without Explanation. An explanation provides comprehensive information that aids in understanding and fixing the vulnerability.
The advent of Large Language Models (LLMs) [11, 12, 13, 14, 15] offers a promising approach to bridging this explanatory gap. A prominent strategy to adapt these general-purpose models for specialized tasks is the “teacher-student” paradigm [16], where a “student” model is fine-tuned on data generated by a powerful “teacher” to distill its capabilities. However, applying this paradigm to the security-critical domain of vulnerability analysis is uniquely challenging. The goal is not merely to generate descriptive text, but to instill a model with the capacity for specialized, autonomous reasoning to both detect and explain vulnerabilities in real-world code. This requires a deep understanding of code semantics and security principles—a level of expertise far beyond general Q&A—presenting a formidable research challenge in transforming LLMs into specialized security assistants, which this work aims to address.
The primary research challenges in this endeavor are twofold: (1) Acquiring High-Quality Vulnerability Explanation Data at Scale: In the field of vulnerability detection, there is a significant lack of annotated explanation data. Additionally, annotating such data typically requires costly manual labeling by domain experts, making the collection of large-scale annotated datasets a major challenge for model fine-tuning. (2) Instilling Autonomous End-to-End Vulnerability Explanation Capability: The goal is to develop a “student” LLM that can autonomously analyze unfamiliar code, detect potential vulnerabilities, and generate accurate, actionable explanations—a task far more demanding than merely explaining pre-identified issues [17]. Effectively fine-tuning a model for this integrated detection-explanation workflow, and determining if learning to explain concurrently enhances detection, remain key open questions.
To address these challenges, we introduce LLMVulExp, a framework centered on a meticulously designed “teacher-student” paradigm that distinguishes between two LLM roles. First, we task a powerful “teacher” LLM with an “open-book” style problem: given the ground-truth vulnerability type and location, it generates a high-quality explanation articulating why the known flaw exists. This rich explanatory data is then used to train a specialized “student” LLM for the “closed-book,” real-world task of autonomously detecting and explaining vulnerabilities in unfamiliar code. Critically, this means the capability distilled from the teacher is not a magical detection ability, but rather the ability to structure and articulate complex security logic, from which the student develops its own autonomous reasoning and detection skills.
To address the first challenge, our approach leverages powerful “teacher” LLMs. By providing them with ground-truth vulnerability type information, we guide them to generate high-quality explanatory data at scale on three vulnerability detection benchmark: SeVC [18], Diversevul [19] and PrimeVul [20]. To ensure rigorous assessment, we also introduce a novel evaluation protocol for the generated explanations, comprising specific metrics—Accuracy, Clarity, and Actionability—and an automated “LLM-as-a-judge” methodology [21, 22], addressing a critical gap in evaluation standards. To address the second challenge, we employ a sophisticated fine-tuning process. Using the synthesized data, we specialize the “student” LLM for its autonomous, dual-capability role: first detecting vulnerabilities in unseen code and then explaining them. Our methodology explicitly investigates whether this explanation-focused training also enhances detection accuracy. We achieve this specialization efficiently using instruction tuning [23] with Low-Rank Adaptation (LoRA) [24] and promote deeper reasoning during inference via a Chain-of-Thought (CoT) prompting strategy [25].
@@
Overall, the primary contributions of this paper are threefold [^1] The replication package of our paper is available [26].:
  • We introduce LLMVulExp, a novel “teacher-student” framework whose core technical innovation is a structured pipeline that leverages guided LLMs to resolve the critical data bottleneck in training explainable vulnerability detection models.
  • We propose and validate a new evaluation protocol for vulnerability explanations, featuring dedicated metrics (Accuracy, Clarity, Actionability) and an “LLM-as-a-judge” method, thereby advancing the assessment standards for LLM-generated security analyses.
  • Through extensive experiments, we demonstrate that our fine-tuning process equips “student” LLMs (e.g., CodeLlama) to autonomously detect vulnerabilities and generate accurate, clear, and actionable explanations. Our results provide valuable insights into the practical application of specialized models for end-to-end vulnerability detection and explanation.
@@ II-A Vulnerability Detection
Prior studies proposed a series of deep learning approaches [1, 2, 6, 7, 8, 27, 28, 29] to detect vulnerabilities. These methods have employed labeled vulnerability data to train neural networks, enabling the models to capture semantic features associated with vulnerabilities.
@@ II-B Vulnerability Explanation
Vulnerability explanation in software security involves providing human-understandable insights into why a piece of code is vulnerable—encompassing its cause, potential exploitation, impact, and remediation—thereby extending beyond mere statement-level localization. Despite its critical importance for aiding developers and reducing reliance on security experts, achieving comprehensive explanations with deep learning models remains a significant, underexplored challenge.
Prior works have made strides in aspects related to Vulnerability explanation. Many, such as VulDeeLocator [38], IVDetect [2], LineVul [39], VELVET [40], and LineVD [41], focus on statement-level vulnerability localization, pinpointing vulnerable code lines using techniques like Bi-LSTMs or graph-based analysis. While this localization is a crucial first step, the outputs often require considerable human interpretation to form a complete explanation. Other studies like VulTeller [10] and VulExplainer [17] offer deeper structural insights that contribute to detection. However, these approaches generally do not generate the complete, natural-language narratives necessary to fully explain the vulnerability’s nature and implications to developers.
Accurately generating comprehensive and human-readable explanations remains a challenge for existing neural models. Our work explores fine-tuning specialized Large Language Models for both vulnerability detection and, crucially, for producing richer, more holistic explanations. We posit that LLMs, with their strong semantic understanding and text generation capabilities, can synthesize information to elucidate the ’why’ and ’how’ of vulnerabilities in coherent language. This aims to move beyond localization or basic structural cues, addressing a critical gap by providing actionable, understandable feedback directly to developers.
@@ III Methodology
❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs requires large quantities of high-quality domain-specific data. In the context of vulnerability detection and explanation, it is crucial to ensure the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples for each type. In this paper, we conduct the study on three C/C++ function-level vulnerability datasets: (1) SeVC [18], which contains four core vulnerability types and over 50,000 vulnerable code snippets, (2) DiverseVul [19], which covers 295 real open-source projects and 150 CWE types and (3) PrimeVul [20], a new vulnerability dataset with high-quality, accurately labeled data and stringent de-duplication to offer realistic training and testing data for vulnerability detection. We select C/C++ as the programming language due to its diverse range of vulnerability types and examples, as well as the substantial dependence of fine-tuning on the availability of large-scale data.
@@
Dataset Non-Vul # Ori. Vul # Ann. Vul # Vul-Type # Eval. Setting
SeVC 364,232 56,395 40,491 4 Single/Multi-Type
DiverseVul 330,492 18,945 9,161 10 Multi-Label
PrimeVul 228,800 6,968 4,041 10 Multi-Class
@@
DiverseVul is a comprehensive C/C++ vulnerability dataset that includes 18,945 vulnerable functions and 330,492 non-vulnerable functions, extracted from 7,514 commits, covering 150 CWEs. As the largest real-world C/C++ vulnerability dataset, DiverseVul is characterized by longer code snippets, a broader range of projects, more diverse vulnerability types, and lower label noise. It serves as a benchmark for evaluating the capability of LLMs in handling real-world scenarios with a wider variety of vulnerabilities. For our fine-tuning and evaluation target, we select the top ten most frequent CWE types, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
PrimeVul is a state-of-the-art benchmark for C/C++ vulnerability detection, designed to overcome the critical data quality issues of prior datasets. It ensures high label accuracy (up to 92% ) through novel automated techniques, such as correlating code changes directly with official NVD descriptions and analyzing single-function security commits. This high-fidelity labeling, combined with stringent de-duplication and realistic chronological data splitting, presents a uniquely challenging testbed. Its difficulty is underscored by the original study, which shows that the performance of powerful models plummets on this benchmark—F1 scores drop from over 68% on older datasets to just 3.09%, and even GPT-4 fails to outperform random guessing on its challenging paired evaluation. Therefore, PrimeVul serves as a stringent benchmark to assess a model’s ability to generalize and reason about complex, real-world vulnerabilities, making it an ideal choice for validating our approach.
We first deduplicate all samples using the SHA-256 hash method and then partition the data into training (80%), validation (10%), and test (10%) sets. To save training cost and mitigate model bias arising from severe class imbalance, we create balanced training and validation sets by downsampling non-vulnerable samples to a 1:1 ratio. The handling of the test set is tailored to the dataset’s characteristics. For the two datasets composed of real-world vulnerabilities, we preserve the original, imbalanced class distribution to ensure the evaluation reflects a realistic deployment scenario. In contrast, for the SeVC dataset, which contains a larger number of synthetic vulnerability samples, we use a balanced test set to facilitate a more effective multi-class evaluation. This overall approach allows for unbiased model training while rigorously testing its practical effectiveness. The final dataset statistics are detailed in Table I.
@@
We implement this annotation process using LLM, selecting appropriate models based on a balance of cost, efficiency, and capability. For our initial annotation of the SeVC and DiverseVul datasets, we employed GPT-3.5 [43], accessed via the API provided by OpenAI [44]. For the subsequently added experiments and the PrimeVul dataset, we utilized Deepseek-V3 [45], a state-of-the-art model noted for its strong code processing capabilities. This dual-model approach allowed us to complete a large-scale annotation effort, resulting in 40,491, 9,161 and 4,041 vulnerability explanation data points across the respective datasets. This effort fills a significant gap in the availability of vulnerability explanation annotations, enabling more effective training and evaluation of vulnerability detection and explanation models. To validate the effectiveness of the synthesized data, we indirectly evaluated the performance of the fine-tuned model in vulnerability detection tasks across multiple scenarios. Additionally, we propose evaluation metrics for vulnerability explanation quality, along with both manual and automated evaluation methods, to directly assess the quality of the generated explanations.
@@ III-B Key Code Extraction Based on Chain-of-Thought (CoT)
To address these challenges and improve the capability of large models to analyze complex code structures, we propose a novel Chain-of-Thought (CoT) enhancement method that incorporates automated key code extraction. This approach enables the model to focus on the most relevant parts of the code, thereby improving accuracy and interpretability. Using the prompt template illustrated in Figure 4, we employ LLMs to automatically extract key statements by analyzing the semantic context of the code and the specific vulnerability type. These extracted statements form a structured CoT, guiding the fine-tuned model through a step-by-step process for targeted vulnerability detection, location, and explanation. For example, in a buffer overflow scenario, key code may include array declarations, bounds-checking statements, or pointer dereferences that contribute directly to the vulnerability. By integrating these extracted key statements into the prompts during fine-tuning, the model can systematically reference this crucial information, enhancing both the interpretability and precision of the vulnerability explanation process.
@@ IV Results
  • RQ4: How does the key code extraction impact detection and explanation performance?
@@ IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
Accurate detection is the foundation for any subsequent vulnerability analysis or explanation. In this RQ, we first conduct a comprehensive evaluation of our fine-tuned model’s detection performance against strong baselines across two distinct, real-world scenarios.
@@
Datasets and Tasks. As detailed in Section III, we select three benchmarks for evaluation: SeVC [18], DiverseVul [19] and PrimeVul [20]. Based on their distinct characteristics, we formulate corresponding detection tasks:
  • On SeVC, we peform binary classification vulnerability toward single-type vulnerability.
  • On SeVC, we perform coarse-grained multi-class vulnerability detection.
  • On DiverseVul, we perform multi-label CWE classification (a function can have multiple vulnerability types) and binary classification.
  • On PrimeVul, we perform multi-class CWE classification (a function has exactly one vulnerability type) and binary classification.
Model and Baselines. Our primary model, LLMVulExp, is based on CodeLlama-13B-Instruct [46], fine-tuned using the LoRA technique. To rigorously situate its performance, we compare it against a comprehensive set of four baselines: two general-purpose pre-trained models, CodeT5 [47] and CodeBERT [48], and two state-of-the-art specialized vulnerability detectors, LineVul [39] and CausalVul [49]. While our ultimate goal includes evaluating explanation capabilities, validating detection efficacy against these strong baselines is a critical first step.
Metrics and Implementation Details. We use Precision, Recall, and F1-Score as our primary evaluation metrics. For the multi-label task on DiverseVul, we report Micro-F1 and Macro-F1. For the multi-class task on SeVC and PrimeVul, we report Weighted-F1 and Macro-F1 to analyze performance across its imbalanced class distribution. We implement our approach using the Transformers [50] and PEFT [51] libraries. All experiments are conducted on two NVIDIA A100-80GB GPUs. Key hyperparameters include a learning rate of 3e-4, 3 training epochs, and a batch size of 2. For the LoRA configuration, we set {q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj}. The LoRA rank is set to 16, the LoRA scaling factor to 16/32, and the dropout rate to 0.05.
@@
IV-A4 RQ1.3: Performance on the DiverseVul Dataset
Task Setup. To evaluate our model’s effectiveness in a standard, realistic scenario, we utilize the DiverseVul dataset. This benchmark is particularly suitable as it reflects real-world conditions where code can contain multiple, distinct types of vulnerabilities. We assess the models on two fronts: (1) a multi-label classification task on the top 10 most frequent CWE types, and (2) a standard binary classification task.
Experimental Results. The comprehensive results on DiverseVul are presented in Table V.
On the binary detection task, a notable observation is that all models achieve a perfect Precision of 100.0%. This suggests the non-vulnerable samples in the test set are sufficiently distinct, leading to a near-zero false positive rate. Therefore, the key differentiating metric is Recall. Here, LLMVulExp demonstrates a decisive advantage, achieving a Recall of 97.7%, which significantly surpasses all baselines by over 34 percentage points. This superiority results in a state-of-the-art F1 Score of 98.8%.
In the more challenging multi-label CWE detection task, our model continues to show robust performance. It achieves the highest Macro-F1 69.9% and Micro-F1 70.5% scores, indicating the most balanced and effective performance overall. This success is primarily driven by its outstanding Macro-Recall and Micro-Recall, once again highlighting its strength in comprehensively identifying all relevant vulnerabilities. While some baselines like CodeT5 attain slightly higher Precision, their significantly lower Recall cripples their overall effectiveness.
These strong results validate the capability of our fine-tuned LLM approach for nuanced vulnerability detection. The example in Figure IV-A4 further illustrates our model’s capacity to generate accurate and context-aware explanations.
TABLE V: Performance Comparison on DiverseVul (RQ1.3). All metric values are percentages (%). Best results in each row are in bold. * means indicates statistically significant improvement over the baseline method through paired t-test ($p<0.05$).
Task Metric Ours CodeBERT CodeT5 LineVul CausalVul
Multi-Label CWE Detection Precision (Micro) 71.7 65.3 72.8 65.0 67.3
Recall (Micro) 69.3 38.2 56.8 38.1 37.4
F1 Score (Micro) 70.5 48.0 63.9 48.0 48.3
Precision (Macro) 72.5 69.0 73.5 68.9 69.0
Recall (Macro) 67.8 32.0 52.7 32.3 32.2
F1 Score (Macro) $\textbf{69.9}^{*}$ 39.0 60.6 40.2 39.1
Binary Vulnerability Detection Precision 100.0 100.0 100.0 100.0 100.0
Recall 97.7 61.1 62.9 63.2 60.7
F1 Score $\textbf{98.8}^{*}$ 75.9 77.2 77.4 75.5
1 Network::FilterStatus Context::onNetworkNewConnection() { 2 onCreate(root_context_id_); 3 if (!wasm_->onNewConnection_) { 4 return Network::FilterStatus::Continue; 5 } 6 if (wasm_->onNewConnection_(this, id_).u64_ == 0) { 7 return Network::FilterStatus::Continue; 8 } 9 return Network::FilterStatus::StopIteration; 10 }